# Orchestrated Robust Controller for the Precision Control of Heavy-duty Hydraulic Manipulators

### v1

@@ Abstract
Vast industrial investment along with increased academic research on hydraulic heavy-duty manipulators has unavoidably paved the way for their automatization, necessitating the design of robust and high-precision controllers. In this study, an orchestrated robust controller is designed to address the mentioned issue. To do so, the entire robotic system is decomposed into subsystems, and a robust controller is designed at each local subsystem by considering unknown model uncertainties, unknown disturbances, and compound input constraints, thanks to virtual decomposition control (VDC). As such, radial basic function neural networks (RBFNNs) are incorporated into VDC to tackle unknown disturbances and uncertainties, resulting in novel decentralized RBFNNs. All these robust local controllers designed at each local subsystem are, then, orchestrated to accomplish high-precision control. In the end, for the first time in the context of VDC, a semi-globally uniformly ultimate boundedness is achieved under the designed controller. The validity of the theoretical results is verified by performing extensive simulations and experiments on a 6-degrees-of-freedom industrial manipulator with a nominal lifting capacity of $600\,kg$ at $5$ meters reach. Comparing the simulation result to state-of-the-art controller along with provided experimental results, demonstrates that the proposed method established all the promises and performed excellently.
@@ I-A Heavy-duty Hydraulic Manipulators
Despite their advantages, HHMs suffer from fundamental challenges, such as model nonlinearities and complexities emerging from structural complicatedness and fluid dynamics[5]. In contrast to electric actuators, whose model complexity can be neglected, hydraulic actuators are governed by fluid dynamics and may be subjected to nonsmooth and discontinuous nonlinearities due to friction between cylinder and piston, compound input constraints, and valve under/overlap. Opposite to the electrically actuated open-chain manipulators, HHMs mostly comprise parallel-serial structures [6] that convert linear piston motion to rotational motion. Meanwhile, in rotary hydraulic actuators, the helical gears are employed to convert the piston motion to rotation, the same mechanism in the wrist actuators of the manipulator in this study (Fig. 1), one of which can deliver $2,000\,N\cdot m$ of torque. The mentioned nonlinearities and complexities render the stability analysis and control design of HHMs significantly challenging, emphasizing the need for robust and high-performance control schemes, not only in theory but also in real-world applications.
@@ I-B State-of-the-Art Review
As explained above, HHMs are characterized by a variety of nonlinearities. Therefore, a nonlinear model-based controller (NMBC) can obtain a better performance than linear controllers[7]. In [8], different NMBC methods, such as sliding mode control (SMC), adaptive inverse dynamics controller (AIDC), and model-reference adaptive controllers with velocity measurement (MRACV), have been designed and compared, unveiling the better performance of MRACV in comparison to AIDC and SMC. In [9], an observer-based adaptive robust controller (ARC) has been designed to control a hydraulic manipulator, in which the backstepping approach has been employed to derive the control law. However, the experimental results in [10] with desired compensation ARC (DCARC) have shown that DCARC yields lower tracking error than ARC. Further, in [11] an adaptive output force tracking controller has been designed to control a 6-degrees-of-freedom (DoF) hydraulic manipulator by employing the virtual decomposition control (VDC) scheme, which was first introduced in [12]. A VDC control, also, in [13] has been designed for the precision control of a hydraulic crane. Additionally, in [14], a VDC-based separate meter-in separate meter-out controller has been presented to not only control the 3-DoF HHM but also improve the energy efficiency of the hydraulic actuator, resulting in a 47% reduction in energy consumption. Moreover, in [15], an adaptive controller with gravity and friction identification has been developed for n-link hydraulic manipulators, and experimental results have been performed for 5-DoF HHM. In [16], a neural adaptive dynamic surface method has been proposed to control the n-link hydraulic manipulator, similar to the traditional backstepping method. The asymptotic convergence of errors with semiglobal boundedness has been proved for the system in the presence of model uncertainty and disturbance. Moreover, the combination of the time delay control with terminal SMC has been utilized to address the control problems of excavators in [17]. A prescribed performance-based adaptive neural controller has been designed in [18] to achieve a good performance for HHM. In both [16] and [18] the radial basic function neural networks (RBFNNs) are employed to estimate uncertainties.
@@
with $|\boldsymbol{e}|_{max}$ and $|\dot{\boldsymbol{x}}_{max}|$ denoting the absolute value of the maximum tracking error and velocity, respectively. A smaller value of $\rho$ demonstrates a better performance for the controller. The $\rho$ value for the controllers in [8], [10], [11],[13], [17], and [14] are 0.0044, 0.005, 0.005, 0.0039, 0.058, and 0.003, respectively. Thus, such a comparison demonstrates that the VDC approach has shown significant performance in comparison to DCARC, SMC, TSMC, and backstepping.
@@
Despite all its benefits, VDC suffers from some drawbacks as well. As expressed in [18] and [16], strong stability in the presence of model uncertainty and disturbance for hydraulic manipulators is both critical and in high demand. However, VDC ensures the $L_{2}$ boundedness for the states of the system which is not strong in the presence of uncertainties, especially when there is compound input constraint, such as deadzone in servo valves and backlash in the gears. Consequently, to have a robust VDC, it is essential to consider unknown model uncertainty, unknown disturbance, and input constraints in the control design while ensuring strong boundedness and stability. These are the open problems in the context of the VDC scheme. Since VDC has shown a better performance than other controllers in the field of HHM, and there is a high demand for HHM in the industry, having a robust VDC would be significantly helpful for paving the way for industrial growth. The contributions of this study properly address the mentioned issues of the VDC.
Due to their universality and perfect capabilities in function approximation, RBFNNs are widely used in robotics and control applications to estimate model uncertainty and disturbances [22, 23]. In [18, 16, 24], uncertainties of hydraulic actuators are estimated using RBFNNs, and uniformly ultimate boundedness (UUB) is proved with asymptotic tracking of the desired trajectory, highlighting the importance of achieving UUB in the presence of uncertainties. In the mentioned works, RBFNN is used to estimate disturbances and uncertainties in a rigid body-actuator model, while separately performing the estimations in actuator and rigid body parts will increase the accuracy. This is especially important in HHMs where the governing dynamics of the actuator and rigid body are completely different (the actuator is governed by fluid dynamics), and separate RBFNNs with corresponding input should be designed to learn the nature of the subsystem. It is stated in [25] that learning for complex systems (like uncertainty and disturbance estimation in 6-DoF HHM) should be solved using the decomposition and re-composition methods. This approach is adopted in our paper for the first time by incorporating RBFNNs into VDC. Decomposing the entire model into subsystems, VDC enables us to design RBFNNs for each subsystem based on the states of that subsystem, like required spatial velocity and acceleration in rigid body subsystems and pressure and piston position and velocity in the actuator subsystem. Then, by orchestrating RBFNNs at the subsystem levels, the re-composed form of estimation is achieved for the complex model. Consequently, the incorporation of RBFNNs into VDC results in novel decentralized RBFNNs (DRBFNNs), feeding which, with the corresponding inputs of a subsystem, will increase the accuracy of the estimation. Such a novel incorporation, despite its mathematical challenges in stability analysis, is important in the sense of performance improvement.
By incorporating input constraints, control systems can operate reliably and prevent undesirable outcomes by respecting the physical limitations of the controlled system or device. The adaptive backlash inverse method introduced in [26] and the adaptive deadzone inverse method introduced in [27] are widely utilized to tackle unknown backlash or deadzone in electric robots [28, 29]. However, a limited amount of research has been conducted to address input constraint in the field of hydraulic manipulators [30, 31, 32]. As stated in [30], backlash and deadzone existing in the valve and gears can result in performance deterioration and even instability. In [33] and [34], the backlash in hydraulic rotary actuators with both rack and pinion mechanism and helical gear has been examined, and a better performance in experiments has been achieved. However, none of the studies considered both backlash and deadzone as compound input constraints in their control method, which is especially crucial for hydraulic actuators because they are more prone to such a compound constraint. In [35], a compound constraint consisting of saturation and backlash/deadzone is converted to an equivalent saturation function using the right inverse of the backlash or deadzone function. However, in real-world applications, the actual parameter of the backlash or deadzone is unavailable, and such an inverse cannot be computed. In this study, the compound input constraint is considered in the control design. The parametrizable part of the deadzone-backlash function is tackled by adaptive control, while the non-parametrizable part is estimated by DRBFNNs to ensure robustness.
I-C Contributions
@@
In this study, compound input constraint, unknown model uncertainty, and unknown disturbance are considered in the VDC scheme to ensure the robustness and perfect performance of the controller. Altogether, this has given rise to the orchestrated robust controller: orchestrating all the local controllers at the rigid body and actuator subsystems that are equipped to achieve high-precision control. As such, the contributions of this study can be stated with their significance as follows:
@@
  • The compound input constraint model is considered in the modeling part and handled by designing a novel adaptive deadzone-backlash inverse controller. The parametrizable part of the deadzone-backlash model is used to handle parameter uncertainties while the non-parametrizable part, which contains all the nonlinearities, is tackled by novel DRBFNNs. In addition, DRBFNNs are employed to estimate unknown model uncertainties and unknown disturbances in actuator and rigid body subsystems.
@@
  • Taking the presented orchestrated robust controller into account along with the complexity of the system, semi-global uniformly ultimate boundedness (SGUUB) is achieved for the first time for VDC. This is a significant contribution, since achieving such stability is mathematically challenging because of the existence of VPFs and the fact that VDC deals with each subsystem.
@@ II Mathematical preliminaries
In this section, the basic mathematical foundation of the VDC approach is presented, which is essential for control design and stability analysis. Moreover, lemmas and assumptions that are employed in the paper for stability analysis are expressed in this section as well. Fig. 1 demonstrates the schematics of the hydraulic robotic system in this study. Fig. 1a shows the kinematic details with base angle ($\zeta_{1}$), lift angle ($\zeta_{2}$), tilt angle ($\zeta_{3}$), and spherical wrist angles of ($\xi_{1}$),($\xi_{2}$), and ($\xi_{3}$). Further, Fig. 1b displays the system that decomposed into three objects, each of which will be analyzed separately, by exploiting the virtual cutting point (VCP) concept of VDC.
II-A VDC Foundation
In the VDC context, linear/angular velocities and forces/moments are properly integrated within a 6D vector to facilitate the transformation of velocities and forces/moments between different frames. Spatial 6D vectors, which are based on Plücker coordinates, allow the complete representation of both linear and angular motion of multibody mechanisms in a compact manner.
Consider $\{A\}$ and $\{B\}$ as frames that are attached to the $i^{th}$ rigid body. Then, the 6D linear/angular velocity vector ${}^{A}V\in\mathbb{R}^{6}$ and force/moment vector ${}^{A}F\in\mathbb{R}^{6}$ can be expressed as follows [12]:
${}^{A}V=[\,^{A}v,\,^{A}\omega]^{T},\quad^{A}F=[\,^{A}f,\,^{A}m]^{T}$
where ${}^{A}v\in\mathbb{R}^{3}$ and ${}^{A}\omega\in\mathbb{R}^{3}$ are the linear and angular velocities of frame $\{A\}$, and ${}^{A}f\in\mathbb{R}^{3}$ and ${}^{A}m\in\mathbb{R}^{3}$ are the force and moment expressed in frame $\{A\}$, respectively. The transformation matrix that transforms force/moment vectors and velocity vectors between frames $\{A\}$ and $\{B\}$ is [12],
@@ II-A VDC Foundation
where ${}^{A}R_{B}\in\mathbb{R}^{3\times 3}$ is a rotation matrix between frame $\{A\}$ and $\{B\}$, and (${}^{A}r_{{A}{B}}\times$) is a skew-symmetric matrix operator defined in [12]. Based on (2), the force/moment and velocity vectors can be transformed between frames as [12],
@@
The net force/moment vector of the rigid body in frame $\left\{A\right\}$, indicated as ${{}^{A}F^{*}}\in\mathbb{R}^{6}$, is:
$*M_{A}\frac{d}{dt}(^{A}V)+C_{A}(^{A}\omega)^{A}V+G_{A}=\,^{A}F^{*}$ (4)
where $M_{A}\in\mathbb{R}^{6\times 6}$ is the mass matrix, $C_{A}\left({{}^{A}\omega}\right)\in\mathbb{R}^{6\times 6}$ is the matrix of Coriolis and centrifugal terms and $G_{A}\in\mathbb{R}^{6}$ includes the gravity terms. Details are expressed in [36].
@@
$M_{A}\frac{d}{dt}\left({{}^{A}V}\right)+C_{A}\left({{}^{A}{\omega}}\right){{}^{A}V}+G_{A}=\,\bar{Y}_{A}\phi_{A}$ (5)
with $\bar{Y}_{A}\in\mathbb{R}^{6\times 10}$ is the compact regressor matrix and $\phi_{A}\in\mathbb{R}^{10}$ is the unique inertial parameter vector.
@@
with $\dot{p}_{d}$ being desired joint velocity, $p_{d}$ being desired joint trajectory, $p$ being measured joint variable, and $\lambda>0$ being a positive constant. The variable $p$ depends on the type of joint; if the joint is revolute, $p$ represents angular measurements, and if the joint is prismatic, $p$ denotes linear motion. Then, the linear/angular velocity vector, indicated as ${{}^{A}V_{r}}$, can be computed based on $\dot{p}_{r}$ by performing kinematic computation, for which the thorough details will be provided later. Thus, the required force/moment vector can be defined as:
@@
where $Y_{A}$ is in the sense of (5) by replacing ${}^{A}V$ with ${}^{A}V_{r}$, $\hat{\phi}_{A}$ is the estimation of $\phi_{A}$, and ${}^{A}F_{c}$ is the regulating control term. The required force/moment vector in (7) demonstrates the amount of force/moment that must be applied to the rigid body subsystem in order to achieve the control objectives at the subsystem level. In the context of VDC, each subsystem (rigid body and actuator) has a corresponding local required control action that accomplishes the local control goals. Then, by orchestrating all these local controllers, the primary goal of the system can be established. Such decomposition into subsystems and orchestration to achieve objectives are due to the VCP and VPF properties of the VDC.
@@
[12]. A non-negative accompanying function $\nu(t)\in{R}$ is a piecewise differentiable function that encompasses the following properties: i) $\nu(t)\leq 0$ for $t>0$, and ii) $\dot{\nu}(t)$ exists almost everywhere.
@@
II-B Lemmas and Assumptions
In this section, we comprehensively enumerate and expound upon all the lemmas and assumptions that serve as foundational elements within the context of this paper.
@@ II-B Lemmas and Assumptions
For an unknown external disturbance $D(t)$ and unknown robot model uncertainties $\Delta_{R}$ and $\Delta_{{a}}$ corresponding to rigid body and actuator subsystems, respectively, we have,
@@
with $\delta_{1},\,\delta_{2},\,\delta_{3}\geq 0$ being unknown constants.

### v2

@@ Abstract
Vast industrial investment along with increased academic research on hydraulic heavy-duty manipulators has unavoidably paved the way for their automatization, necessitating the design of robust and high-precision controllers. In this study, an orchestrated robust controller is designed to address the mentioned issue for a generic manipulator with an anthropomorphic arm and spherical wrist. To do so, the entire robotic system is decomposed into subsystems, and a robust controller is designed at each local subsystem by considering unknown model uncertainties, unknown disturbances, and compound input nonlinearities, thanks to virtual decomposition control (VDC). As such, radial basic function neural networks (RBFNNs) are incorporated into VDC to tackle unknown disturbances and uncertainties, resulting in novel decentralized RBFNNs. All robust local controllers designed at each local subsystem, then, are orchestrated to accomplish high-precision control. In the end, for the first time in the context of VDC, a semi-globally uniformly ultimate boundedness is achieved under the designed controller. The validity of the theoretical results is verified by performing extensive simulations and experiments on a 6-degrees-of-freedom industrial manipulator with a nominal lifting capacity of $600\,kg$ at $5$ meters reach. Comparing the simulation result to the state-of-the-art controller along with provided experimental results, demonstrates that proposed method established all promises and performed excellently.
Index Terms:
Hydraulic manipulators, neuro-adaptive control, input nonlinearities, nonlinear model-based control
List of Abbreviations
HHM Heavy-duty hydraulic manipulator
PWR Power-to-weight ratio
DoF Degree of freedom
VDC Virtual decomposition control
ORC Orchestrated robust controller
DRBFNN Decentralized radial basis function neural network
UUB Uniformely ultimate boundedness
SGUUB semi-globall UUB
VPF Virtual power flow
VCP Virtual cutting point
RHA Rotary hydraulic actuator
NMBC nonlinear model-based controller
List of Symbols
$\xi_{i},\zeta_{i}\in\mathbb{R}$ Joint angles
{A}, {B} Body frame of a rigid body
${}^{A}V\in\mathbb{R}^{6}$ Spatial velocity vector
${}^{A}v\in\mathbb{R}^{3}$ linear velocity vector
${}^{A}\omega\in\mathbb{R}^{3}$ angular velocity vector
${}^{A}F\in\mathbb{R}^{6}$ Spatial force vector
${}^{A}f\in\mathbb{R}^{3}$ force vector
${}^{A}\tau\in\mathbb{R}^{3}$ moment vector
${}^{A}F^{*}\in\mathbb{R}^{6}$ Net spatial force vector
${}^{A}V_{r}\in\mathbb{R}^{6}$ Required spatial velocity vector
${}^{A}F_{r}\in\mathbb{R}^{6}$ Required spatial force vector
${}^{A}F^{*}_{r}\in\mathbb{R}^{6}$ Net required spatial force vector
${}^{A}U_{B}\in\mathbb{R}^{6\times 6}$ Transformation matrix
${}^{A}R_{B}\in\mathbb{R}^{3\times 3}$ Rotation matrix
${}^{A}r_{{A}{B}}\in\mathbb{R}^{3}$ Distance between {A} and {B}
$M_{A}\in\mathbb{R}^{6\times 6}$ Mass matrix
$C_{A}\in\mathbb{R}^{6\times 6}$ Coriolis and centrifugal matrix
$G_{A}\in\mathbb{R}^{6}$ Gravitational forces vector
$Y_{A}\in\mathbb{R}^{6\times 10}$ Regressor matrix based on ${}^{A}V_{r}$
$\phi_{A}\in\mathbb{R}^{10}$ Inertial parameter vector
$p_{A}\in\mathbb{R}$ Virtual power flow (VPF)
$\nu(t)$ Non-negative accompanying function
$\dot{\nu}(t)$ Time derivative of $\nu(t)$
$\delta_{i},\,i=1,2,3$ Unknown positive contants
$\Delta_{R}(t)\in\mathbb{R}^{6}$ Rigid body model uncertainty
$\Delta_{a}(t)\in\mathbb{R}$ Actuator model uncertainty
$D(t)\in\mathbb{R}^{6}$ Unknown disturbance
$\Psi(.)$ Radial basis function (RBF)
$\chi$ Input vector of the activation function
$W^{*}$ Optimal weights of neural networks (NNs)
$Z(.)$ Output of RBFNNs
$\mathcal{L}_{A}\in\mathbb{R}^{4\times 4}$ Pseudo-inertia matrix
$h\in\mathbb{R}^{3}$ first mass moment
$\bar{I}\in\mathbb{R}^{6\times 6}$ Rotational inertia matrix
$m\in\mathbb{R}$ mass
$x_{(.)}(t),\dot{x}_{(.)}(t)$ position and velocity of the piston
$r_{p},r_{wi}$ gear ratio
$q_{j}(t),q_{j1}(t),q_{j2}(t)$ closed-chain angles
$\dot{q}_{j}(t),\dot{q}_{j1}(t),\dot{q}_{j2}(t)$ closed-chain velocities
$L_{j},L_{j1},L_{j2}$ Link lengths
$x_{j0}$ actuator effective length
$f_{(.)}(t)$ Linear hydraulic actuator force
$J(t)\in\mathbb{R}^{6\times 6}$ Jacobian matrix
$\dot{X}_{d}(t),\dot{Y}_{d}(t),\dot{Z}_{d}(t)$ End-effector desired linear velocity
$\dot{\varpi}_{xd}(t),\,\dot{\varpi}_{yd}(t),\,\dot{\varpi}_{zd}(t)$ End-effector desired angular velocity
$\dot{\Pi}_{d}(t)\in\mathbb{R}^{6}$ End-effector desired velocity vector
$\dot{\Theta}_{d}(t)\in\mathbb{R}^{6}$ Joint desired velocity vector
$\lambda,\bar{\lambda}_{xj},\sigma_{i}$ Positive constants
$K_{A},\Gamma$ Positive-definite constants
$u$ Spool valve voltage
$DZ(.)$ Deadzone nonlienarity
$BS(.)$ Backlash nonlinearity
$DB(.)$ Deadzone-backlash nonlinearity
$m_{d}>0,b_{r}>0,b_{l}<0$ Deadzone parameters
$v^{*}$ Deadzone input
$u^{*},u^{*}_{d}$ Backlash and desired backlash outputs, respectively
$k_{b}>0,B_{r}>0,B_{l}<0$ Backlash parameters
$\bar{\hat{w}}(t)$ Estimation of discontinuous function
$\theta\in\mathbb{R}^{5}$ Deadzone-backlash parameter vector
$\eta\in\mathbb{R}^{5}$ Deadzone-backlash regression vector
$\eta_{0}$ Deadzone-backlash mismatch error
$Y_{f}\in\mathbb{R}^{1\times 7}$ Friction regressor matrix
$\phi_{f}\in\mathbb{R}^{7}$ Friction parameter vector
$A_{a},A_{b}$ Cross-sectional area
$p_{a}(t),p_{b}(t)$ Chamber pressures
$p_{s}$ Supply pressure
$p_{r}$ Oil tank pressure
$c_{p1},c_{p2},c_{n1},c_{n2}$ Flow coefficients
$Q_{a}(t),Q_{b}(t)$ Flow rate
$s$ Maximum stroke of piston
$c_{l}$ Leakage coefficient
$\beta$ oil bulk modulus
$\theta_{v},\theta_{d}$ Actuator parameter vectors
$Y_{v}(t),Y_{d}(t)$ Regressor matrices
$k_{f},k_{x}$ Positive constants
$\gamma,\pi,\gamma_{(.)},\gamma_{(.)}$ Positive constants
$\hat{\square}$ Estimated signals
$\tilde{\square}$ Estimation error ($\square-\hat{\square}$)
@@ I-A Heavy-duty Hydraulic Manipulators
Despite their advantages, HHMs suffer from fundamental challenges, such as model nonlinearities and complexities emerging from structural complicatedness and fluid dynamics[5]. In contrast to electric actuators, whose model complexity can be neglected, hydraulic actuators are governed by fluid dynamics and may be subjected to nonsmooth and discontinuous nonlinearities due to friction between cylinder and piston, compound input nonlinearities, and valve under/overlap. Opposite to the electrically actuated open-chain manipulators, HHMs mostly comprise parallel-serial structures [6] that convert linear piston motion to rotational motion. Meanwhile, in rotary hydraulic actuators, the helical gears are employed to convert the piston motion to rotation, the same mechanism in the wrist actuators of the manipulator in this study (Fig. 1), one of which can deliver $2,000\,N\cdot m$ of torque. The mentioned nonlinearities and complexities render the stability analysis and control design of HHMs significantly challenging, emphasizing the need for robust and high-performance control schemes, not only in theory but also in real-world applications.
@@ I-B State-of-the-Art Review
As explained above, HHMs are characterized by a variety of nonlinearities. Therefore, NMBC can obtain a better performance than linear controllers[7]. In the following, the proposed schemes for HHMs control objectives are elaborated on.
I-B1 Model-based controller
In [8], different NMBC methods, such as sliding mode control (SMC), adaptive inverse dynamics controller (AIDC), and model-reference adaptive controllers with velocity measurement (MRACV), have been designed and compared, unveiling the better performance of MRACV in comparison to AIDC and SMC. In [9], an observer-based adaptive robust controller (ARC) has been designed to control a hydraulic manipulator, in which the backstepping approach has been employed to derive the control law. However, the experimental results in [10] with desired compensation ARC (DCARC) have shown that DCARC yields lower tracking error than ARC. Further, in [11] an adaptive output force tracking controller has been designed to control a 6-degrees-of-freedom (DoF) hydraulic manipulator by employing the virtual decomposition control (VDC) scheme, which was first introduced in [12]. A VDC control, also, in [13] has been designed for the precision control of a hydraulic crane. Additionally, in [14], a VDC-based separate meter-in separate meter-out controller has been presented to not only control the 3-DoF HHM but also improve the energy efficiency of the hydraulic actuator, resulting in a 47% reduction in energy consumption. In [15], an adaptive controller with gravity and friction identification has been developed for n-link hydraulic manipulators, and experimental results have been performed for 5-DoF HHM. Moreover, the combination of the time delay control with terminal SMC has been utilized to address the control problems of excavators in [16]. In both [17] and [18] the radial basic function neural networks (RBFNNs) are employed to estimate uncertainties.
@@ I-B1 Model-based controller
with $|\boldsymbol{e}|_{max}$ and $|\dot{\boldsymbol{x}}_{max}|$ denoting the absolute value of the maximum tracking error and velocity, respectively. A smaller value of $\rho$ demonstrates a better performance for the controller. The $\rho$ value for the controllers in [8], [10], [11],[13], [16], and [14] are 0.0044, 0.005, 0.005, 0.0039, 0.058, and 0.003, respectively. Thus, such a comparison demonstrates that the VDC approach has shown significant performance in comparison to DCARC, SMC, TSMC, and backstepping. Consequently, the VDC scheme has been employed as the baseline controller in the current study.
I-B2 Input nonlinearities
By incorporating input nonlinearities, control systems can operate reliably and prevent undesirable outcomes by respecting the physical limitations of the controlled system or device. The adaptive backlash inverse method introduced in [19] and the adaptive deadzone inverse method introduced in [20] are widely utilized to tackle unknown backlash or deadzone in electric robots [21, 22]. However, a limited amount of research has been conducted to address input constraint in the field of hydraulic manipulators [23, 24, 25]. As stated in [23], backlash and deadzone existing in the valve and gears can result in performance deterioration and even instability. In [26] and [27], the backlash in hydraulic rotary actuators with both rack and pinion mechanism and helical gear has been examined, and a better performance in experiments has been achieved. However, none of the studies considered both backlash and deadzone as compound input nonlinearities in their control method, which is especially crucial for hydraulic actuators because they are more prone to such a compound constraint. In [28], a compound constraint consisting of saturation and backlash/deadzone is converted to an equivalent saturation function using the right inverse of the backlash or deadzone function. However, in real-world applications, the actual parameter of the backlash or deadzone is unavailable, and such an inverse cannot be computed. Therefore, considering the compound input nonlinearities in the control law is an open problem and needs to be addressed.
I-B3 Unmodeled uncertainties
Due to their universality and perfect capabilities in function approximation, RBFNNs are widely used in robotics and control applications to estimate model uncertainty and disturbances [29, 30]. In [17], a neural adaptive dynamic surface method has been proposed to control the n-link hydraulic manipulator, similar to the traditional backstepping method. The asymptotic convergence of errors with SGUUB has been proved for the system in the presence of model uncertainty and disturbance. A prescribed performance-based adaptive neural controller has been designed in [18] to achieve a good performance for HHM. In [31], uncertainties of hydraulic actuators are estimated using RBFNNs and UUB is proved with asymptotic tracking of the desired trajectory, highlighting the importance of achieving UUB in the presence of uncertainties. In the mentioned works, RBFNN is used to estimate disturbances and uncertainties in a rigid body-actuator model, while separately performing the estimations in actuator and rigid body parts will increase the accuracy. This is especially important in HHMs where the governing dynamics of the actuator and rigid body are completely different (the actuator is governed by fluid dynamics), and separate RBFNNs with corresponding input should be designed to learn the nature of the subsystem. It is stated in [32] that learning for complex systems (like uncertainty and disturbance estimation in 6-DoF HHM) should be solved using the decomposition and re-composition methods. Consequently, having a decentralized RBFNNs to deal with actuator and rigid body dynamics can boost the control performance and is of high importance.
I-B4 VDC scheme
@@
Despite all its benefits, VDC suffers from some drawbacks as well. As expressed in [18] and [17], strong stability in the presence of model uncertainty and disturbance for hydraulic manipulators is both critical and in high demand. However, VDC ensures the $L_{2}$ boundedness for the states of the system which is not strong in the presence of uncertainties, especially when there is compound input nonlinearities, such as deadzone in servo valves and backlash in the gears. Consequently, to have a robust VDC, it is essential to consider unknown model uncertainty, unknown disturbance, and input nonlinearities in the control design while ensuring strong boundedness and stability. These are the open problems in the context of the VDC scheme. Since VDC has shown a better performance than other controllers in the field of HHM, and there is a high demand for HHM in the industry, having a robust VDC would be significantly helpful for paving the way for industrial growth. The contributions of this study properly address the mentioned issues of the VDC.
I-C Aim and Contributions
I-C1 Aim
The aim of the current paper is to achieve a robust and high precision control law for a generic 6-DoF HHMs consisting of anthropomorphic arm and spherical wrist. For doing so, the ORC scheme is proposed to consider compound input nonlinearities along with unmodeled uncertainties to achieve a robust and accurate tracking performance.
I-C2 Contributions
@@
Considering all the above-mentioned issues, compound input nonlinearities, unknown model uncertainties, and unknown disturbances are incoporated into control law to ensure the robustness and perfect performance of the controller. Altogether, this has given rise to the ORC scheme: orchestrating all the local controllers at the rigid body and actuator subsystems that are equipped to achieve high-precision control. As such, the contributions of this study can be stated with their significance as follows:
@@
  • The compound input nonlinearities model is considered in the modeling part and handled by designing a novel adaptive deadzone-backlash inverse controller. The parametrizable part of the deadzone-backlash model is used to handle parameter uncertainties while the non-parametrizable part, which contains all the nonlinearities, is tackled by novel DRBFNNs. In addition, DRBFNNs are employed to estimate unknown model uncertainties and unknown disturbances in actuator and rigid body subsystems.
@@
  • Taking the presented ORC into account along with the complexity of the system, semi-global uniformly ultimate boundedness (SGUUB) is achieved for the first time for VDC. This is a significant contribution, since achieving such stability is mathematically challenging because of the existence of VPFs and the fact that VDC deals with each subsystem.
@@ II Mathematical preliminaries
In this section, the basic mathematical foundation of the VDC approach along with important lemmas and assumptions are presented, which are essential for control design and stability analysis. Fig. 1 demonstrates the schematics of the generic 6-DoF HHM with anothromorphic arm and spherical wrist. Fig. 1a shows the kinematic details with base angle ($\zeta_{1}$), lift angle ($\zeta_{2}$), tilt angle ($\zeta_{3}$), and spherical wrist angles of ($\xi_{1}$),($\xi_{2}$), and ($\xi_{3}$). Further, Fig. 1b displays the system that decomposed into three objects, each of which will be analyzed separately, by exploiting the virtual cutting point (VCP) concept of VDC.
In the VDC context, spatial velocities and forces/moments are properly integrated within a 6D vector to facilitate the transformation of velocities and forces/moments between different frames. Spatial 6D vectors, which are based on Plücker coordinates, allow the complete representation of both linear and angular motion of multibody mechanisms in a compact manner.
Consider $\{A\}$ and $\{B\}$ as frames that are attached to the $i^{th}$ rigid body. Then, the spatial velocity and force vectors can be expressed as follows [12]:
${}^{A}V=[\,^{A}v,\,^{A}\omega]^{T},\quad^{A}F=[\,^{A}f,\,^{A}\tau]^{T}$
The transformation matrix that transforms spatial force and velocity vectors between frames $\{A\}$ and $\{B\}$ is [12],
@@
where ($\times$) is a skew-symmetric operator defined in [12]. Based on (2), the spatial force and velocity vectors can be transformed between frames as [12],
@@
The net spatial force vector of the rigid body in frame $\left\{A\right\}$ is [36]:
$M_{A}\frac{d}{dt}(^{A}V)+C_{A}(^{A}\omega)^{A}V+G_{A}=\,^{A}F^{*}.$ (4)
@@
$M_{A}\frac{d}{dt}\left({{}^{A}V}\right)+C_{A}\left({{}^{A}{\omega}}\right){{}^{A}V}+G_{A}=\,\bar{Y}_{A}\phi_{A}.$ (5)
@@
with $\dot{p}_{d}$ being desired joint velocity, $p_{d}$ being desired joint trajectory, $p$ being measured joint variable, and $\lambda>0$ being a positive constant. The variable $p$ depends on the type of joint; if the joint is revolute, $p$ represents angular measurements, and if the joint is prismatic, $p$ denotes linear motion. Then, the spatial velocity vector, indicated as ${{}^{A}V_{r}}$, can be computed based on $\dot{p}_{r}$ by performing kinematic computation, for which the thorough details will be provided later. Thus, the required spatial force vector can be defined as:
@@
where $Y_{A}$ is in the sense of (5) and ${}^{A}F_{c}$ is the regulating control term. The required spatial force vector in (7) demonstrates the amount of spatial force that must be applied to the rigid body subsystem in order to achieve the control objectives at the subsystem level. In the context of VDC, each subsystem (rigid body and actuator) has a corresponding local required control action that accomplishes the local control goals. Then, by orchestrating all these local controllers, the primary goal of the system can be established. Such decomposition into subsystems and orchestration to achieve objectives are due to the VCP and VPF properties of the VDC.
@@
[12]. A non-negative accompanying function $\nu(t)\in{R}$ is a piecewise differentiable function that encompasses the following properties: i) $\nu(t)\geq 0$ for $t>0$, and ii) $\dot{\nu}(t)$ exists almost everywhere.
Definition 4
[12]. A decomposed subsystem is called virtually stable if for the given $\nu(t)$, the $\dot{\nu}(t)\leq-\|B(t)\|+VPFs$ can be established, with $B(t)$ being the positive function.
@@
For the unknown signals, we have,
@@
with $\delta_{1},\,\delta_{2},\,\delta_{3}\geq 0$ being unknown constants, indicating the upper bounds.